Give the domain of the function.

6a. g(x)=x^3-1

6b. f(t)= 10/(t^2-2t-3)

6c. h(x)= square root of 2x+1
1
Expert's answer
2012-12-12T09:17:01-0500
6a. x in (-infinity, +infinity)
6b.find the roots of denominator
t^2-2t-3 = 0
t1=(2+sqrt(4+3*4))/2= 3
t2=(2-sqrt(4+3*4))/2= -1
hence domain is (-infinity, +infinity) except x=3 andx=-1
6c. sqrt(2x+1)
2x+1>0
domain is x>-1/2